50cm3 of carbon (11) oxide was exploded with 150cm3 of air containing 20% oxygen by volume, which of the reactants was in excess?
1
Expert's answer
2018-08-13T02:25:46-0400
V(CO) = 50 cm^3 = 50 ml = 0.05 l
V(air) = 150 cm^3 = 150 ml = 0.15 l
V(O2)/V(air) = 20%
_____________________________
2 CO + O2 = 2CO2

n(CO) = V(CO)/V(M) = 0.05/22.4 = 0.0022mol
V(O2)/V(air) = x/150 x=150×0.2 =30 ml=0.03l
n(O2) = V(O2)/V(M) =0.03/22.4 = 0.0013 mol
From the reaction ratio n(CO): n(O2) = 2:1
For explosing 0.0022 mol CO need 0.0011 mol O2. We have 0.0013 mol O2, then , 0.0003 mol O2 is excess.
Answer: O2 was in excess